OpenStudy (anonymous):

find the volume obtained when shaded region is rotated through 360 about the x-axis

OpenStudy (anonymous):



OpenStudy (amistre64):

the area of a circle is pi r^2 the sum of the areas of the circles relates to the integration of the areas\[\int pi~r^2\] \[\pi \int ~r^2\] since the radius is defined by f(x) \[\pi \int_{a}^{b} ~(f(x))^2~dx\]

OpenStudy (anonymous):

Should i replace x = secx..?

OpenStudy (amistre64):

the radius is defined by f(x) = sec(x) yes so integrating r^2 = sec^2(x) is appropriate

OpenStudy (anonymous):

thx buddy.. :)

OpenStudy (amistre64):

youre welcome